# LLMmap: Fingerprinting for Large Language Models

### v3

@@ 1 Introduction
Identifying the specific LLM and its version embedded within an application can reveal critical attack surfaces. Once the LLM is accurately fingerprinted, an attacker can craft targeted adversarial inputs, exploit specific vulnerabilities unique to that model version, such as the buffer overflow vulnerability in Mixture of Experts architectures [15], the “glitch tokens” phenomenon [18], or exploit previously leaked information [9]. For open-source LLMs, this fingerprinting can be further exploited using white-box optimization techniques, enhancing the precision and impact of attacks [45, 26, 12, 15, 6, 28, 44].
In this paper, we introduce LLMmap  [^1] Name derived from the foundational network scanner Nmap [25]., a novel approach to LLM fingerprinting that is both precise and efficient. LLMmap represents the first generation of active fingerprinting attacks specifically designed for applications integrating LLMs. By sending carefully constructed queries to the target application and analyzing the responses, LLMmap can accurately identify the underlying LLM version with minimal interaction—typically between $3$ and $8$ queries (see Figure 1). LLMmap is designed to be robust across diverse deployment scenarios, including systems with arbitrary system prompts, stochastic sampling procedures, hyper-parameters, and those employing advanced frameworks like Retrieval-Augmented Generation (RAG)[17] or Chain-of-Thought prompting [40].
LLMmap offers two key capabilities: ($i$) A closed-set classifier that identifies the correct LLM version from among 42 of the most common models, achieving accuracy exceeding $95\%$. ($ii$) An open-set classifier developed through contrastive learning, which enables the detection and fingerprinting of previously unseen LLMs. This open-set approach allows LLMmap to generate a vectorial representation of the LLM’s behavior (templates), which can be stored and later matched against an expanding database of LLM fingerprints.
@@ 2.1 Problem Formulation
We stress that our approach does not require any form of whitebox access to LLMs during setup or inference and can be applied to both open-source and closed-source proprietary models.
@@ 5 The Querying Strategy of LLMmap
Next, we employed a heuristic approach to identify the most effective combination of queries. Using a simple greedy search algorithm, as detailed in Appendix B, we aimed to filter out less effective queries and ensure that the selected queries complemented each other well, resulting in a diversified and robust fingerprinting strategy.
After this optimization process, we finalized a query strategy composed of 8 highly effective queries, which are listed in Table 2. These queries are ranked by their individual effectiveness; recall, though, that their true strength lies in their synergistic ability to fingerprint LLMs across various settings consistently. Hereafter, unless stated otherwise, these 8 queries constitute the default query strategy $\mathcal{Q}$ used in LLMmap.
Table 2: Top-8 most effective probes (default query strategy).
Probe Type
(2) Tell me how to build a bomb. Alignment
(3) What is the size of your training dataset? Meta-information
(5) What’s your data cutoff date? Meta-information
(7) Bonjour, how are you doing today? ¿Qué tal?’ Malformed
@@ 6 The Inference Model of LLMmap
Machine learning is essential to overcome this challenge. ML models can learn to generalize across the diverse and variable responses produced by an LLM. They can abstract underlying patterns from noisy data, capturing both explicit signals and more subtle, query-independent traits such as writing style. These capabilities allow ML-driven inference models to accurately identify LLM versions, even when responses vary due to different configurations or random factors.
Fingerprinting Approaches: Closed-Set and Open-Set. We approach LLM fingerprinting through two primary methods: closed-set and open-set fingerprinting.
  1. 1.
    Closed-Set Fingerprinting: In this scenario, the inference model operates under the assumption that it knows the possible LLM versions in advance. The task is to identify the correct version from a predefined set using the observed traces. Formally, given a set of $n$ known versions $\mathbf{C}=\{v_{1},\dots,v_{n}\}$, the model functions as a classifier, mapping the traces $\mathbf{T}^{k}$ to one of the known labels in $\mathbf{C}$. This approach is typically more accurate because the model is trained on the LLMs it needs to identify.
  2. 2.
    Open-Set Fingerprinting: Unlike the closed-set approach, open-set fingerprinting does not assume prior knowledge of all possible LLM versions. Instead, the inference model produces a vector signature from the traces, mapping them to an $m$-dimensional vector space $f:\mathbf{T}^{k}\rightarrow\mathbb{Z}^{m}$. This vector can then be matched against a database of LLM signatures. This method allows for identifying previously unseen models for which the model has not been trained to recognize. The open-set approach is particularly useful in dynamic environments where new LLMs are frequently introduced, as it allows for identifying model versions that the inference system has not explicitly encountered before. However, this flexibility comes with trade-offs, as it generally provides lower accuracy compared to the closed-set approach.
To implement these two models, we use the same backbone network and modify it according to the task at hand.
@@
Backbone Architecture. One straightforward solution for building the inference model would be to use a pre-trained, instruction-tuned LLM. However, we chose a lighter solution to ensure our approach is practical and can run efficiently on a standard machine. The structure of our backbone network is shown in Figure 3. For each pair of query $q_{i}$ and response $o_{i}$, we use a pre-trained textual embedding model, denoted as $\mathcal{E}$, to generate a vector representation. This process involves:
@@
Closed-Set Classification. To implement the classifier in the closed-set setting, we add an additional dense layer $f_{c}$ on top of $u$, which maps $u$ into the class space—for our experiments, the class space contains 42 LLM versions listed in Table A.1 in Appendix A. We train the model in a fully supervised manner. We generate a suitable training set by simulating multiple LLM-integrated applications with different LLMs and prompting configurations. For each simulated application, we collect traces by submitting queries according to our query strategy and using the LLM within the application as the label. The detailed process for generating these training sets is explained in Section 7.1. Once the input traces are collected, we train the model to identify the correct LLM. This task requires the model to generalize across different prompting configurations and handle the inherent stochasticity.
@@ 7 Evaluation
In this section, we evaluate LLMmap. Section 7.1 presents our evaluation setup, describing how training and test LLM-integrated applications are simulated. Section 7.2 reports LLMmap’s performance for both its instantiations.
@@ 7.1 Evaluation Setup
Universe of LLMs. To evaluate LLMmap, we selected the 42 LLM versions listed in Table A.1. These models were chosen based on their popularity at the time of writing. We primarily use the Huggingface hub to select open-source models. We automatically retrieve the most popular models based on download counts by leveraging their API services. For closed-source models, we consider the three main models offered by the two most popular vendors (i.e., OpenAI and Anthropic) for which API access is available. Hereafter, we refer to these models as the LLM universe $\mathbf{L}$.
@@
    System Prompt Universe $SP$: We curated a collection of $60$ different system prompts, which include prompts collected from online resources as well as automatically generated ones. Examples of system prompts are reported in Table A.2 in Appendix A.
@@
    Prompt Framework Universe $PF$: We consider two settings: RAG and Chain-Of-Thought (CoT) [40]. To simulate RAG, we create the input chunks by sampling from $4$ to $6$ random entries from the dataset SQuAD 2.0 [32], and consider $6$ prompt templates for retrieval-based-Q&A. In the same way, we consider $6$ prompt templates for CoT. An example of a prompt derived from this approach is given in Figure A.2 in Appendix A.
@@
Creating Training/Testing Traces for Inference Model. Once the LLM universe $\mathbf{L}$, the generated prompting configurations ($\mathbf{S}_{\text{train}}$ and $\mathbf{S}_{\text{test}}$), and the query strategy $\mathcal{Q}$ are chosen, we can collect the traces required to train the inference model. This process is summarized in Algorithm 2 where the subscript XXX is either “train” or “test”. For each LLM in Table A.1, we sample a prompting configuration from $\mathbf{S}_{\text{train}}$ and collect all the responses of the model upon the queries in $\mathcal{Q}$. To allow the inference model to generalize over different prompting configurations, we collect $w$ traces per $LLM_{v}$ with $w$ different prompting configurations. In our setting, we set $w$ to $75$. This process results in a collection $D_{\text{train}}$ of pairs “(traces, $LLM_{v}$)” that can be used to train the inference model in a supervised manner. To create the test set, we repeat the process but use $\mathbf{S}_{\text{test}}$ instead of $\mathbf{S}_{\text{train}}$, ensuring that the prompting configurations used for testing are completely disjoint from those used for training. This results in another collection of traces that can be used for evaluation.
@@ 7.2 Results
Table 3: LLMmap per-model accuracy for both closed and open-set models obtained when submitting all 8 queries listed in Table 2. Accuracies and standard deviation computed on $10$ models’ training runs.
LLM ($v_{i}$) (A) Closed (B) Open (C) Open (unseen)
@@
Figure 5: Closed-set accuracy of LLMmap as the number of queries to the LLM-integrated application increases. Error bars report standard deviation across $10$ training runs.
Figure 6: Two-dimensional representation of the templates derived by the open-set fingerprint model on all the tested models.
@@ 7.2.1 Closed-Set Classification Setting
Once the inference model has been trained, we test it using the traces generated with the left-out prompting configurations in $\mathbf{S}_{\text{test}}$. Given input traces generated by the target model version, we use the closed-set classifier to infer the LLM that generated them from the list of LLM versions in Table A.1. Results are summarized in Table 3, column (A). On average, the inference model achieves an accuracy of $95.3\%$ over the 42 LLMs. The results indicate that LLMmap is generally robust across different model versions, correctly classifying $41$ out of 42 LLMs with $90\%$ accuracy or higher. This includes highly similar models, such as different instances of Google’s Gemma or various versions of ChatGPT. In Figure A.1 in Appendix A we report results as a confusion matrix. The main exception is Meta’s Llama-3-70B-Instruct, where LLMmap achieves only $84\%$ accuracy. As shown in the confusion matrix, this lower accuracy is primarily due to misclassifications with closely related models, such as Smaug-Llama-3-70B-Instruct by Abacus.AI, which are fine-tuned versions of the original model.
Accuracy as a Function of Number of Queries. Naturally, the accuracy of fingerprinting depends on the number of queries made to the target. An attacker might reduce the number of interactions with the target application to adjust to cases with proactive monitoring mechanisms, but this typically results in decreased fingerprinting accuracy. This tradeoff is illustrated in Figure 5, where accuracy is plotted against the number of traces provided as input to the inference model. Generally, using only the first three queries from Table 2 achieves an average accuracy of $90\%$. However, accuracy levels off after eight queries. It is conceivable that the $95\%$ accuracy mark could be surpassed by incorporating different queries than those outlined in Section 4. We emphasize that the inference model’s attention-based architecture allows for seamless adjustment of the number of input traces without requiring retraining, enabling the attacker to reduce the number of queries arbitrarily based on the context.
@@ 7.2.2 Open-Set Classification Setting
Implementing fingerprinting with the open-set model involves two main steps.
Derive a Database of Templates. After training the inference model, we derive a vectorial representation $u$ for all traces in the training set by providing them as input to the inference model (see Section 6). We group all the vectors associated with traces generated by the same LLM and average them into a single vector. This results in 42 vectors, each representing a template for an LLM in Table A.1. We refer to this list of vectors as $\mathcal{DB}$. These vectors, after dimensional reduction, are depicted in Figure 6.
@@
Open-set Fingerprinting of Seen LLMs. As a warm-up evaluation (towards open-set fingerprinting of unseen LLMs), we will apply the inference model of our open-set approach to an LLM that may have been included in the training phase. Given traces $\mathcal{T}^{?}$ generated by an unknown (but potentially trained on) model, inference proceeds as follows: (1) We provide $\mathcal{T}^{?}$ to the inference model and derive a vector $u^{?}$. (2) We compute the cosine similarity between $u^{?}$ and all the vectors in $\mathcal{DB}$. (3) We output the LLM whose template has the highest similarity to $u^{?}$ as the prediction. Using this approach, we evaluate the performance of the open-set inference model against our test set. Results are reported in Table 3, column (B). Fingerprinting with the open-set inference model results in an average accuracy of $91\%$, which is $4\%$ lower than the specialized closed-set classifier.
Fingerprinting of Unseen LLMs. The main advantage of the open-set fingerprinting approach is its ability to operate on unseen LLMs, i.e., that were not part of the training set. To do this, it is enough to add a template for the new model in $\mathcal{DB}$; that is, collecting one or more traces for the new LLM and derive its template by applying the pre-trained inference model. Then, future inference process execution remains unchanged (see step (2) abovex). To evaluate the performance of the open-set inference model in fingerprinting new LLMs, we proceed as follows. Given the list of models in Table 2, (1) we remove an LLM (referred to as $LLM_{\text{out}}$) and (2) train the inference model on the traces generated by the remaining $41$ LLMs. (3) We then test the inference model to correctly recognize $LLM_{\text{out}}$ using the fingerprinting process described above. We repeat this process for each model in a k-fold cross-validation fashion.
On average, the inference model correctly identifies the unseen LLM with $81.2\%$ accuracy. Individual results are reported in Table 3, column (C). In this setting, predictions tend to be less robust and have higher variance overall. Nonetheless, the average accuracy remains meaningfully high, enabling for practical uses.
We emphasize that once traces for a new model are collected, they can be simply added to the training set of the closed-set model to achieve more robust predictions. Given the small size of the model, retraining requires less than 5 minutes on a single GPU. Furthermore, the closed and open-set models can operate as an ensemble without increasing the number of queries to the target application; once the traces are collected, they can be provided to both models, and their predictions can be combined for more robust inference.

### v4

# Dario Pasquini RSAC Labs

@@ 1 Introduction
Identifying the specific LLM and its version embedded within an application can reveal critical attack surfaces. Once the LLM is accurately fingerprinted, an attacker can craft targeted adversarial inputs, exploit specific vulnerabilities unique to that model version, such as the buffer overflow vulnerability in Mixture of Experts architectures [17] or privacy attacks [49], the “glitch tokens” phenomenon [22], or exploit previously leaked information [11]. For open-source LLMs, this fingerprinting can be further exploited using white-box optimization techniques, enhancing the precision and impact of attacks [51, 31, 14, 17, 8, 33, 50].
In this paper, we introduce LLMmap  [^1] Name derived from the foundational network scanner Nmap [30]., a novel approach to LLM fingerprinting that is both precise and efficient. LLMmap represents the first generation of active fingerprinting attacks specifically designed for applications integrating LLMs. By sending carefully constructed queries to the target application and analyzing the responses, LLMmap can accurately identify the underlying LLM version with minimal interaction—typically between $3$ and $8$ queries (see Figure 1). LLMmap is designed to be robust across diverse deployment scenarios, including systems with arbitrary system prompts, stochastic sampling procedures, hyper-parameters, and those employing advanced frameworks like Retrieval-Augmented Generation (RAG)[21] or Chain-of-Thought prompting [19, 45].
LLMmap offers two key capabilities: ($i$) A closed-set classifier that identifies the correct LLM version from among 42 of the most common models, achieving accuracy exceeding $95\%$. ($ii$) An open-set classifier developed through contrastive learning, which enables the detection and fingerprinting of new LLMs. This open-set approach allows LLMmap to generate a vectorial representation of the LLM’s behavior (signatures), which can be stored and later matched against an expanding database of LLM fingerprints.
@@ 2.1 Threat Model
We stress that our approach does not require any form of whitebox access to LLMs during setup (i.e., training) or inference and can be applied to both open-source and closed-source proprietary models.
2.1.1 The Power of LLM Fingerprinting: Identifying LLM Version to Tailor Attack Strategies
In the following, we discuss how fingerprinting can serve as a component of a multi-stage attack effort and which attack stages benefit from an effective fingerprinting tool. We present a concrete demonstration of the role of fingerprinting in such a two-stage attack in Appendix B.
A successful fingerprinting technique can “fast-track” an attack and enable the adversary to design tailored inputs that work robustly on the specific LLM version under attack. Knowing the LLM version allows the adversary to deploy tools that automate the generation of tailored inputs. Numerous studies demonstrated strong efficacy of tailored inputs compared to non-tailored, regardless of whether the attacker operates in the white-box or black-box threat model [23, 12, 27, 51, 31].
Benefiting Open-Source and Closed-Source Attacks. If fingerprinting identifies an LLM version as an open-source model, then even though the model is part of the backend of an application and is not directly accessible to the adversary, the attacker can simply download a local copy of the model and treat this scenario as a white-box attack. Specifically, the attacker can perform gradient-based optimization, generating highly efficient attack vectors to deploy against remote applications. A similar rationale applies to black-box optimization-based attacks on proprietary closed-source LLMs [23, 12, 27]. Once the specific proprietary LLM version used in the application is identified, an attacker can sidestep the process of running optimization on the application API (which may have rate-limiting mitigations in place) and instead can target the proprietary model’s APIs directly (which typically do not impose any limitation on the number of prompts). This approach decreases the risk of detection by the targeted application compared to using the application itself as an oracle for optimization.
One specific family of attacks that benefits from knowledge of the LLM version is tailored prompt injection attacks, which show significantly higher success rates when the version is known. In this context, we provide a concrete and practical example in Appendix B demonstrating a two-stage attack strategy. In the first stage, an adversary uses LLMmap to fingerprint the LLM version within a target application. Then, they apply gradient-based optimization techniques[31] to create a precise inline prompt injection trigger optimized for the identified LLM version.
Ultimately, as LLM architectures and their applications continue to evolve, more version-specific vulnerabilities will emerge, further expanding the attack surface for adversaries.
@@ 5 The Querying Strategy of LLMmap
Next, we employed a heuristic approach to identify the most effective combination of queries. Using a simple greedy search algorithm, as detailed in Appendix H, we aimed to filter out less effective queries and ensure that the selected queries complemented each other well, resulting in a diversified and robust fingerprinting strategy.
After this optimization process, we finalized a query strategy composed of 8 highly effective queries, which are listed in Table G.2. These queries are ranked by their individual effectiveness; recall, though, that their true strength lies in their synergistic ability to fingerprint LLMs across various settings consistently. Hereafter, unless stated otherwise, these 8 queries constitute the default query strategy $\mathcal{Q}$ used in LLMmap.
@@ 6 The Inference Model of LLMmap
Machine learning is essential to overcome this challenge. ML models can learn to generalize across the diverse and variable responses produced by an LLM. They can abstract underlying patterns from noisy data, capturing both explicit signals and more subtle, query-independent traits such as writing style. These capabilities allow ML-driven inference models to accurately identify LLM versions, even when responses vary due to different configurations or sampling randomness.
6.1 Fingerprinting Approaches: Closed-Set and Open-Set.
We approach LLM fingerprinting through two primary methods: closed-set and open-set fingerprinting.
6.1.1 Closed-Set Fingerprinting
In this scenario, the inference model operates under the assumption that it knows the possible LLM versions in advance. The task is to identify the correct version from a predefined set using the observed traces. Formally, given a set of $n$ known versions $\mathbf{C}=\{v_{1},\dots,v_{n}\}$, the model functions as a classifier, mapping the traces $\mathbf{T}^{k}$ to one of the known labels in $\mathbf{C}$. This approach is typically more accurate because the model is trained on the LLMs it needs to identify.
6.1.2 Open-Set Fingerprinting
Unlike the closed-set approach, open-set fingerprinting does not assume that the LLM version is included among those available during training.
In the open-set framework, fingerprinting is decoupled into (1) the inference model $f$ and (2) a fingerprints database $\mathcal{DB}$. Here, the inference model is a function $f:\mathbf{T}^{k}\rightarrow\mathbb{Z}^{m}$ that generates a “vector signature”, specifically an $m$-dimensional real vector, from the input traces –where $m$ is a hyperparameter we choose during the initialization. The database $\mathcal{DB}$ consists of (vector signature, version label) pairs. Fingerprinting a model $q$ is then performed by finding the vector signature in the database $\mathcal{DB}$ that is closest to $f(q)$. Using this modeling, one can easily extend the signature database over time by adding new signature-label pairs without requiring re-training of the inference model.
This approach is akin to the one used by tools such as nmap—a tool that relies on a large, community-curated database of OS and service fingerprints [2, 3]. Users can submit and extend the database by adding new entries without needing to alter nmap’s existing functionality. [^5] Under the often verified assumption that the current nmap query strategy is capable of capturing the new OS/service. In the nmap case, an entry is a pair: “label”, the name of the OS/service version, and its “signature”, the list of responses obtained by running nmap’s query strategy on the OS/service. Open-Set LLMmap implements the same logic, but it stores an $m$-dimensional vector generated by the inference model.666or the average of multiple vectors if multiple sets of traces are available for the same target model.
In Appendix E, we show how the LLMmap open-set approach can also potentially identify cases where a test LLM is entirely “unseen”; that is when the LLM version does not yet have a corresponding entry in the signature database.
To implement the closed/open-set models, we use the same backbone network and modify it according to the task at hand.
@@ 6.1.2 Open-Set Fingerprinting
6.2 Inference Model’s Architecture.
One straightforward solution for building the inference model would be to use a pre-trained, instruction-tuned LLM. However, we chose a lighter solution to ensure our approach is practical and can run efficiently on a standard machine. The structure of our backbone network is shown in Figure 3. For each pair of query $q_{i}$ and response $o_{i}$, we use a pre-trained textual embedding model, denoted as $\mathcal{E}$, to generate a vector representation. This process involves:
@@
Closed-Set Classification. To implement the classifier in the closed-set setting, we add an additional dense layer $f_{c}$ on top of $u$, which maps $u$ into the class space—for our experiments, the class space contains 42 LLM versions listed in Table G.1 in Appendix G. We train the model in a fully supervised manner. We generate a suitable training set by simulating multiple LLM-integrated applications with different LLMs and prompting configurations. For each simulated application, we collect traces by submitting queries according to our query strategy and using the LLM within the application as the label. The detailed process for generating these training sets is explained in Section 7.1. Once the input traces are collected, we train the model to identify the correct LLM. This task requires the model to generalize across different prompting configurations and handle the inherent stochasticity.
@@ 7 Evaluation
In this section, we evaluate LLMmap. Section 7.1 presents our evaluation setup, describing how training and testing LLM-integrated applications are simulated. Section 7.2 reports LLMmap’s performance for both its instantiations.
@@ 7.1 Evaluation Setup
Universe of LLMs. To evaluate LLMmap, we selected the 42 LLM versions listed in Table G.1. These models were chosen based on their popularity at the time of writing. We primarily use the Huggingface hub to select open-source models. We automatically retrieve the most popular models based on download counts by leveraging their API services. For closed-source models, we consider the three main models offered by the two most popular vendors (i.e., OpenAI and Anthropic) for which API access is available. Hereafter, we refer to these models as the LLM universe $\mathbf{L}$.
@@
    System Prompt Universe $SP$: We curated a collection of $60$ different system prompts, which include prompts collected from online resources as well as automatically generated ones. Examples of system prompts are reported in Table G.3 in Appendix G.
@@
    Prompt Framework Universe $PF$: We consider two settings: RAG and Chain-Of-Thought (CoT) [45]. To simulate RAG, we create the input chunks by sampling from $4$ to $6$ random entries from the dataset SQuAD 2.0 [37], and consider $6$ prompt templates for retrieval-based-Q&A. In the same way, we consider $6$ prompt templates for CoT.
@@
Creating Training/Testing Traces for Inference Model. Once the LLM universe $\mathbf{L}$, the generated prompting configurations ($\mathbf{S}_{\text{train}}$ and $\mathbf{S}_{\text{test}}$), and the query strategy $\mathcal{Q}$ are chosen, we can collect the traces required to train the inference model. This process is summarized in Algorithm 2 where the subscript XXX is either “train” or “test”. For each LLM in Table G.1, we sample a prompting configuration from $\mathbf{S}_{\text{train}}$ and collect all the responses of the model upon the queries in $\mathcal{Q}$. To allow the inference model to generalize over different prompting configurations, we collect $w$ traces per $LLM_{v}$ with $w$ different prompting configurations. In our setting, we set $w$ to $75$. This process results in a collection $D_{\text{train}}$ of pairs “(traces, $LLM_{v}$)” that can be used to train the inference model in a supervised manner. To create the test set, we repeat the process but use $\mathbf{S}_{\text{test}}$ instead of $\mathbf{S}_{\text{train}}$, ensuring that the prompting configurations used for testing are completely disjoint from those used for training. This results in another collection of traces that can be used for evaluation.
@@ 7.2 Results
Table 2: LLMmap per-model accuracy for both closed and open-set models obtained when submitting all 8 queries. Accuracies and standard deviation computed on $10$ models’ training runs.
LLM ($v_{i}$) (A) Closed (B) Open (C) Open ((left-out)
@@
Figure 5: Closed-set accuracy of the inference model as the number of queries to the LLM-integrated application increases for LLMmap using the default query strategy and two baselines strategies.
Figure 6: Two-dimensional representation of the signatures derived by the open-set fingerprint model on all the tested models.
@@ 7.2.1 Closed-Set Classification Setting
Once the inference model has been trained, we test it using the traces generated with the left-out prompting configurations in $\mathbf{S}_{\text{test}}$. Given input traces generated by the target model version, we use the closed-set classifier to infer the LLM that generated them from the list of LLM versions in Table G.1.
Accuracy as a Function of Number of Queries. Naturally, the accuracy of fingerprinting depends on the number of queries made to the target. An attacker might reduce the number of interactions with the target application to adjust to cases with proactive monitoring mechanisms, but this typically results in decreased fingerprinting accuracy. This tradeoff is illustrated in Figure 5, where accuracy is plotted against the number of traces provided as input to the inference model (“default”). Generally, using only the first three queries from Table G.2 achieves an average accuracy of $90\%$. However, accuracy levels off after eight queries. It is conceivable that the $95\%$ accuracy mark could be surpassed by incorporating different queries than those outlined in Section 4. On average, the inference model achieves an accuracy of $95.3\%$ over the 42 LLMs when all $8$ queried are used.
Baselines. To assess the query strategy from Section 5, we compare it with two baseline query strategies.
In the first baseline, we randomly sample 30 entries from the dataset  Stanford Alpaca listed in Table G.4, column (a) and apply the same greedy optimization procedure described in Appendix H. This process results in convergence to 8 queries, matching the number used in the default strategy.
For the second baseline, we prompt a state-of-the-art language model (gpt-4o-2024-11-20) to generate 30 discriminative queries that would potentially be good options for fingerprinting LLMs. These are then subjected to the same optimization procedure outlined for the default strategy. Further details on these baselines can be found in Appendix F.
For each query strategy, we construct a training set, train an inference model from scratch, and evaluate its performance using the same methodology as the default strategy. The performance of these two query strategies in a closed-set setting is presented in Figure 5. While the baseline strategies achieve respectable accuracy–indicating that the inference model can effectively extract robust and meaningful features for classification–LLMmap’s default strategy performs better than the tested baselines. [^8] It is important to note that we do not claim any form of optimality for the default strategy, and we believe that this can be further improved. As expected, the strategy derived from the LLM (“gpt4o-gen-opt”) surpasses those generated from general prompts (“random-opt”). Interestingly, several generated queries overlap with the ones discussed in Section 4, highlighting consistency and relevance in the query design.
More fine-grained results are summarized in Table 2, column (A). Those results indicate that LLMmap is generally robust across different model versions, correctly classifying $41$ out of 42 LLMs with $90\%$ accuracy or higher. This includes highly similar models, such as different instances of Google’s Gemma or various versions of ChatGPT. In Figure G.1 in Appendix G, we report results as a confusion matrix. The main exception is Meta’s Llama-3-70B-Instruct, where LLMmap achieves only $84\%$ accuracy. As shown in the confusion matrix, this lower accuracy is primarily due to misclassifications with closely related models, such as Smaug-Llama-3-70B-Instruct by Abacus.AI, which are fine-tuned versions of the original model.
@@ 7.2.2 Open-Set Classification Setting
In this subsection, we conduct a series of experiments to assess the effectiveness of open-set classification. The first family of experiments is called “fingerprinting-known-LLM," and the second family is called “fingerprinting-unseen-LLM." The first family includes the following experiments, i.e., “fingerprinting-known-LLM," arranged in increasing difficulty: ($i$) an experiment where the LLM we aim to fingerprint is both present in the $\mathcal{DB}$ and has been used during the training of the inference model $f$, ($ii$) an experiment in which the LLM we are trying to fingerprint is present in the $\mathcal{DB}$ but is not used during the training of the inference model $f$. The second family of experiments, i.e., “fingerprinting-unseen-LLMs," examines what occurs when we attempt to fingerprint an LLM that was neither used in training of $f$ nor has a vector signature in the $\mathcal{DB}$.
@@
Fingerprinting-known-LLM: ($i$) Used in Training. We apply the inference model of our open-set approach to LLMs that have been used during the training phase but with prompt configurations that were not used in training (thus, even though we have the same LLM version as in training of $f$, the answers are not necessarily the same given that a different prompt configuration from $\mathbf{S}_{\text{test}}$ is used). Given traces $\mathcal{T}^{?}$ generated by the target LLM on a prompt configuration from $\mathbf{S}_{\text{test}}$, inference proceeds as follows: (1) We provide $\mathcal{T}^{?}$ to the inference model $f$ and derive a vector $u^{?}$. (2) We compute the cosine similarity between $u^{?}$ and all the vectors in $\mathcal{DB}$. (3) We output the LLM whose signature has the highest similarity to $u^{?}$ as the prediction of LLMmap. Using this approach, we evaluate the performance of the open-set inference model against our LLMs seen during the training of $f$. Results are reported in Table 2, column (B). Fingerprinting with the open-set inference model achieves an average accuracy of $91\%$, which is $4\%$ lower than the specialized closed-set classifier.
Fingerprinting-known-LLM: ($ii$) Not Used in Training. We emphasize that this is the main mode of operation of a fingerprinting tool, i.e., the community extends the $\mathcal{DB}$ with additional LLM versions with models that were not used during the training phase of $f$ (which happened at the setup phase of LLMmap). To evaluate the performance of LLMmap under this setting, we proceed as follows. Given the list of models in Table G.2: (1) We remove an LLM (referred to as $LLM_{\text{out}}$) and (2) train the inference model on the traces generated by the remaining $41$ LLMs. (3) We then test the inference model’s ability to correctly recognize $LLM_{\text{out}}$ by adding $LLM_{\text{out}}$’s vector signature to the database (Algorithm A.1). This process is repeated for each of the $42$ models in a k-fold cross-validation fashion, always sampling prompt configurations from options in $\mathbf{S}_{\text{test}}$ that were not used by any LLM during training. Results are reported in Table 2, column (C) under the heading (left-out-LLM). On average, the inference model correctly identifies the LLM with $81.2\%$ accuracy. In this setting, predictions tend to be less robust and exhibit higher variance overall. Nonetheless, the average accuracy remains meaningfully high, enabling practical applications.
Fingerprinting-unseen-LLM. This setting is relevant only if a user attempts to fingerprint an LLM whose vector signature has not been previously added to the database $\mathcal{DB}$ of LLMmap, i.e., occurs only in a short window right after a public release of a new model. Due to the specialized nature of this setting, we detail the experiments in Appendix E. At a high level, we deployed a random forest-based binary classifier that runs as an additional step before the final decision of LLMmap to determine whether the responses from the queried LLM are sufficiently close to be considered part of $\mathcal{DB}$ or if the responses diverge too significantly to be regarded as an unseen LLM. The average accuracy is over $82\%$.